# Data-Efficient Operator Learning via Unsupervised Pretraining and In-Context Learning

### v1

@@ Appendix A PDEs
We consider two time-independent PDEs (Poisson, Helmholtz) and two time-dependent PDEs (Reaction-Diffusion, Navier-Stokes), as described below.
@@
    ${\nabla}u+\omega u=f\quad\text{in}\ \Omega,$ (2)
@@
    where $\{D_{u},D_{v}\}$ are diffusion coefficients for $u$ and $v$, respectively, and where $R_{u}$ and $R_{v}$ are reaction functions, which are defined as the Fitzhugh-Nagumo equation [Klaasen and Troy, 1984]:
@@ Appendix B Detailed Experiment Settings
B.1 Unlabeled PDE Data.
  1. 1.
    For Poisson and Helmholtz [Li et al., 2021a, Subramanian et al., 2023], we consider coefficients, sourcing functions, and coordinates as unlabeled PDE data. These data are controlled by the physical parameters of PDE systems (diffusion coefficients, wavenumber).
  2. 2.
    For the Diffusion-Reaction and Navier Stokes problems, we follow PINO [Li et al., 2021b] and PDEBench [Takamoto et al., 2022] to use snapshots without temporal dynamics as the unlabeled PDE data, controlled by diffusion coefficients and Reynolds number, respectively.
B.2 Distributions of Unlabeled PDE Data.
In Table 1, we summarize the distribution of our unlabeled PDE data during pretraining, fine-tuning, and OOD inference with ICL. Ranges of these physical parameters are inspired by Subramanian et al. [2023]. During pretraining, we consider a wide distribution of unlabeled PDE data. When training (fine-tuning) our model, we consider in-distribution unlabeled PDE data. Finally, we test our ICL method with OOD settings.333 For Navier Stokes from PDEBench, we use the original data. We could not generate our own pretraining/finetuning data with different ranges of physics parameters, due to a possible mismatch of the provided configuration files and the version of Phiflow used in PDEBench (see GitHub issue at https://github.com/pdebench/PDEBench/issues/36). For Helmholtz OOD, we choose a narrow range of coefficients ([15, 20]) mainly because its solution is very sensitive to the wavenumber, and FNO’s performance significantly drops when we move to more extreme OOD settings.
@@
B.3 Hyperparameters
@@
We summarize our hyperparameters used during pretraining and fine-tuning/training in Table 2. These hyperparameters strictly follow previous works [Li et al., 2021a, Subramanian et al., 2023, Takamoto et al., 2022, Li et al., 2021b, McCabe et al., 2023].
@@
Table 2: Hyperparameters for pretraining and training/fine-tuning. “N.S.”: 2D Incompressible Navier-Stokes. “DAdapt”: adaptive learning rate by D-adaptation [Defazio and Mishchenko, 2023]. “ns”: total number of simulated training samples.
@@
Stage $\downarrow$ PDEs $\rightarrow$ Poisson Helmholtz Reaction-Diffusion N.S. (PINO) N.S. (PDEBench)
Pretraining Number of Samples 46,080 46,080 76,760 57,545 713,286
Learning Rate $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ DAdapt
Batch Size 32 32 5 2 8
Resolution ($H\times W$) 64$\times$64 64$\times$64 128$\times$128 128$\times$128 512$\times$512
Epochs/Iterations 500 epochs 500 epochs 500 epochs 100,000 iters 500 epochs
Training/Fine-tuning Learning Rate $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ DAdapt
Batch Size $\min$(32, ns) $\min$(32, ns) 5 2 8
Resolution ($H\times W$ or $H\times W\times T$) 64$\times$64 64$\times$64 128$\times$128$\times$10 64$\times$64$\times$1 512$\times$512$\times$15
Epochs/Iterations 500 epochs 500 epochs 500 epochs 50,000 iters 500 epochs
Rollouts N/A N/A 91 N/A 1
Appendix C Ablation Studies
C.1 Magnitude of Perturbations during Pretraining.
@@ Appendix C Ablation Studies
C.2 Ablation of the Number of Pretraining Samples.
@@
Appendix D Visualization of MAE Pretraining
MAE Pretraining on Different PDEs.
@@ Appendix D Visualization of MAE Pretraining
Figure 6: Visualization of reconstructions of unlabeled PDE data on the Poisson (“Pois.”), Helmholtz (“Helm.”), 2D Diffusion-Reaction (“D.R.”), and 2D incompressible Navier-Stokes (“N.S.”) equations during MAE pretraining. (Mask ratio: 0.1 for Poisson, Helmholtz, and 2D Diffusion-Reaction equations; 0.7 for incompressible Navier-Stokes.) In masks, only white areas are visible to the model during pretraining.
Figure 7: Visualization of reconstructions of unlabeled PDE data on the 2D incompressible Navier-Stokes equations during MAE pertaining with mask ratio from 0.1 to 0.9.
Appendix E Example of Simulation Costs
In Table 8, we demonstrate the cheap simulation of only unlabeled PDE data, versus simulating both unlabeled PDE data and solutions, on 2D incompressible Navier-Stoke on PINO Dataset [Li et al., 2021b] and Reaction-Diffusion on PDE-Bench [Takamoto et al., 2022]. We can see that unlabeled PDE data are extremely cheap to simulate. Therefore, our pretraining method can boost the performance and meanwhile save the heavy cost of data simulations.
Table 8: Simulation time costs on 2D Incompressible Navier-Stokes (“N.S.”) on PINO Dataset and Reaction-Diffusion (“R.D.”) on PDE-Bench. “$Re$”: Reynolds number. “$D_{u},D_{v}$”: diffusion coefficients.
Data Physical Parameters Unlabeled PDE Data (s) Data+Solutions (s) Data Size CPU GPU
N.S. $Re=100$ 5499.32 11013.90 $20\times 100\times 21\times 512\times 512$ 1 AMD EPYC 7763 1 NVIDIA A100 (40GB)
$Re=300$ 3683.02 7625.82
$Re=500$ 4059.71 8963.39
$Re=800$ 4829.3 10811.15
$Re=1000$ 4957.24 10788.69
R.D. $D_{u}=1\times 10^{-3},D_{v}=5\times 10^{-3}$ 29.65 6657.34 $1000\times 101\times 128\times 128\times 2$ 1 AMD EPYC 7763 N/A
Appendix F Benefits of In-context Learning (ICL)

### v2

@@ Appendix A PDEs
We consider two time-independent PDEs (Poisson, Helmholtz) and two time-dependent PDEs (Reaction-Diffusion, Navier-Stokes), as described below. We also describe types of unlabeled data for per-PDE examples. For a general definition of unlabeled PDE data, please refer to Section 3.1.1.
@@
    Unlabeled PDE Data: In [11], inputs to neural operators have four channels: the source function, and three diffusion coefficients (two diagonal elements and one off-diagonal element in the symmetric diffusion coefficient tensor) expanded to the whole spatial domain. We use this input as the unlabeled data to pretrain the neural operator.
@@
    $u+\omega u=f\quad\text{in}\ \Omega,$ (2)
@@
    Unlabeled PDE Data: In [11], inputs to neural operators have two channels: the source function, and the wavenumber $\omega$ expanded to the whole spatial domain. We use this input as the unlabeled data to pretrain the neural operator.
@@
    where $\{D_{u},D_{v}\}$ are diffusion coefficients for $u$ and $v$, respectively, and where $R_{u}$ and $R_{v}$ are reaction functions, which are defined as the Fitzhugh-Nagumo equation [82]:
@@
    Unlabeled PDE Data: In PDEBench [10], we forecast the activator $u$ and the inhibitor $v$ (i.e. two input channels) with $T=10$. Since each individual snapshot can serve as an initial condition for forecasting, during our unsupervised pretraining we discard the temporal dynamics and use randomly shuffled snapshots of $u$ and $v$ with $T=1$ as unlabeled PDE data. That means, during pretraining the model only has access to single and individual snapshots without any awareness of temporal dynamics.
@@
    Unlabeled PDE Data: We follow the training settings from previous works. 1) For FNO [4], we focus on mapping the vorticity up to time $T$ to the vorticity at a later time $T+1$ ($\left.w_{[0,T]}\mapsto w\right|_{T+1}$) (with $T=33$). The input includes four channels: the vorticity, and three (xy-spatial and temporal) linear mesh position embedding channels. Thus, similar to Reaction-Diffusion above, during pretraining, the model will take individual snapshots of vorticity and position embeddings as unlabeled data. 2) For PDEBench [10], we forecast both xy velocities and vorticity (i.e. three input channels), with $T=15$. The model will take individual snapshots of vorticity and xy velocities as unlabeled data.
@@ Appendix B Detailed Experiment Settings
B.1 Distributions of Unlabeled PDE Data
In Table 1, we summarize the distribution of our unlabeled PDE data during pretraining, fine-tuning, and OOD inference with ICL. Ranges of these physical parameters are inspired by [11]. During pretraining, we consider a wide distribution of unlabeled PDE data. When training (fine-tuning) our model, we consider in-distribution unlabeled PDE data. Finally, we test our ICL method with OOD settings.333 For Navier Stokes from PDEBench, we use the original data. We could not generate our own pretraining/finetuning data with different ranges of physics parameters, due to a possible mismatch of the provided configuration files and the version of Phiflow used in PDEBench (see GitHub issue at https://github.com/pdebench/PDEBench/issues/36). For Helmholtz OOD, we choose a narrow range of coefficients ([15, 20]) mainly because its solution is very sensitive to the wavenumber, and FNO’s performance significantly drops when we move to more extreme OOD settings.
@@ B.1 Distributions of Unlabeled PDE Data
B.2 Data Generation
Unlabeled PDE Data.
We generate data for Poisson and Helmholtz [11], Reaction-Diffusion on PDE-Bench [10] and 2D incompressible Navier-Stokes on PINO Dataset [9] following the procedure mentioned in the paper. For unlabeled data generation, we bypass the computation of solvers, which expedites the generation speed, as shown in Table 3.
OOD Samples.
The OOD data generation procedure is similar to the unlabeled data, except for the changes in the physical parameters coefficients. For Poisson and Helmholtz, we consider changing the range of diffusion eigenvalue and waver number respectively. For Navier-Stokes equation, we change the Reynolds number. We list the coefficients in Table 1.
B.3 Training Hyperparameters
@@
We summarize our hyperparameters used during pretraining and fine-tuning/training in Table 2. These hyperparameters strictly follow previous works [4, 11, 10, 9, 47]. We conducted our experiments on four A100 GPUs, each with 40GB of memory.
@@
Table 2: Hyperparameters for pretraining and training/fine-tuning. “N.S.”: 2D Incompressible Navier-Stokes. “DAdapt”: adaptive learning rate by D-adaptation [83]. “ns”: total number of simulated training samples. A batch size of “$\min$(32, ns)” is because the total number of training samples might be fewer than 32.
@@
Stage $\downarrow$ PDEs $\rightarrow$ Poisson Helmholtz Reaction-Diffusion N.S. (PINO) N.S. (PDEBench) ERA5 ScalarFlow Airfoil
Pretraining Number of Samples 46,080 46,080 76,760 57,545 713,286 8,760 56,160 43,104
Learning Rate $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ DAdapt $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Batch Size 32 32 5 2 8 32 32 32
Resolution: ($H\times W$) 64$\times$64 64$\times$64 128$\times$128 128$\times$128 512$\times$512 180$\times$180 180$\times$120 96$\times$45
Epochs/Iterations 500 epochs 500 epochs 500 epochs 100,000 iters 500 epochs 500 epochs 500 epochs 500 epochs
Training / Fine-tuning Learning Rate $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ DAdapt $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Batch Size $\min$(32, ns) $\min$(32, ns) 5 2 8 4 4 $\min$(32, ns)
Resolution:
($H\times W$ or $H\times W\times T$)
64$\times$64 64$\times$64 128$\times$128$\times$10 64$\times$64$\times$1 512$\times$512$\times$15 180$\times$180$\times$16 180$\times$120$\times$16 96$\times$45
Epochs/Iterations 500 epochs 500 epochs 500 epochs 50,000 iters 500 epochs 500 epochs 500 epochs 500 epochs
Rollouts N/A N/A 91 N/A 1 N/A N/A N/A
Appendix C Example of Simulation Costs
In Table 3, we demonstrate the cheap simulation of only unlabeled PDE data, versus simulating both unlabeled PDE data and solutions, on 2D incompressible Navier-Stoke on PINO Dataset [9] and Reaction-Diffusion on PDE-Bench [10]. We can see that unlabeled PDE data are extremely cheap to simulate. Therefore, our pretraining method can boost the performance and meanwhile save the heavy cost of data simulations.
Table 3: Simulation time costs on 2D Incompressible Navier-Stokes (“N.S.”) on PINO Dataset [9] and Reaction-Diffusion (“R.D.”) on PDE-Bench [10]. “$Re$”: Reynolds number. “$D_{u},D_{v}$”: diffusion coefficients. $N$: number of samples. $T$: temporal resolution. $H\times W$: spatial resolution. $C$: input channels (1 for the vorticity in N.S., 2 for velocities $u,v$ in R.D.).
Data
Physical
Parameters
Unlabeled PDE Data
Data+Solutions
Data Size CPU GPU
N.S. $Re=100$ 5499.32 11013.90 $N\times T\times H\times W=$ $20\times 2000\times 512\times 512$ 1 AMD EPYC 7763 1 NVIDIA A100 (40GB)
$Re=300$ 3683.02 7625.82
$Re=500$ 4059.71 8963.39
$Re=800$ 4829.3 10811.15
$Re=1000$ 4957.24 10788.69
R.D.
$D_{u}=1\times 10^{-3}$
$D_{v}=5\times 10^{-3}$
29.65 6657.34
$N\times T\times H\times W\times C=$
$1000\times 101\times 128\times 128\times 2$
1 AMD EPYC 7763 N/A
Appendix D Model Architectures
In Figure 7, We show visualizations of architectures described in Sec. 3.1.4.
Figure 7: Visualizations of architectures we studied. Left: FNO [9]. Right: VideoMAE [65].
Appendix E Comparison with Contrastive Learning
Contrastive learning is an important self-supervised pretraining technique studied in computer vision. For a fair comparison, we directly compare with MoCo v2 [16], a highly-cited self-supervised learning method also originally implemented in PyTorch, whose core method is closely related to SimCLR [14] (originally implemented in TensorFlow).
We compare MoCo v2 with our method on a broader real-world benchmark ERA5 [78]. As shown in Figure 8, our unsupervised learning method can largely outperform MoCo v2. This extra comprehensive result demonstrates that our method can be widely adopted in real-world problems, outperforming previous unsupervised learning methods.
Figure 8: Comparison between our unsupervised pretraining method versus MoCo v2 [16].
Appendix F Consistently Improved Generalization
Beyond the generation gap we have shown in Figure 4 (left) for the Navier Stokes equation from PDEBench, we further collect generation gaps of our models learned on other PDEs. As shown in Figure 9, on diverse PDE systems, our method can contribute to universally reduced overfitting.
Figure 9: Universally reduced overfitting (i.e. smaller generalization gaps) on diverse PDEs (a to d).
Appendix G More Ablation Studies
G.1 Magnitude of Perturbations during Pretraining.
@@ Appendix G More Ablation Studies
G.2 Ablation of the Number of Pretraining Samples.
@@
Figure 10: Visualization of FNO reconstructions of unlabeled PDE data on the Poisson (“Pois.”), Helmholtz (“Helm.”), 2D Diffusion-Reaction (“D.R.”), and 2D incompressible Navier-Stokes (“N.S.”) equations during MAE pretraining. (Mask ratio: 0.1 for Poisson, Helmholtz, and 2D Diffusion-Reaction equations; 0.7 for incompressible Navier-Stokes.) In masks, only white areas are visible to the model during pretraining.
Appendix H Visualization of MAE Pretraining
@@
Figure 11: Visualization of FNO reconstructions of unlabeled PDE data on the 2D incompressible Navier-Stokes equations during MAE pertaining with mask ratio from 0.1 to 0.9.
Appendix I Benefits of In-context Learning (ICL)